Please Help!
"the graph of x+3y-2z=20 contains which point?"
A. (0,10,5)
B.(2,6,2)
C.(6,8,5)
D.(0,0,-10)
I had A as my answer and it was wrong.
Thanks so much!

Answer by Alan3354(30937)

"the graph of x+3y-2z=20 contains which point?"
A. (0,10,5)
B.(2,6,2)
C.(6,8,5)
D.(0,0,-10)
I had A as my answer and it was wrong.
----------------
x+3y-2z=20
A. (0,10,5)
0 + 3*10 - 2*5 =? 20
30 - 10 = 20 so A is on the plane.
----------------
B. (2,6,2)
2 + 3*6 - 2*2 =? 20
16 <> 20 B is not
------------------
C. (6,8,5)
6 + 3*8 - 2*5 =? 20
30 - 10 = 20 C is on the plane
------------
Check D